### 19 Mar 2024 v1

# Transformer Neural Opeartors

Abstract
Existing neural operator architectures face challenges when solving multiphysics problems with coupled partial differential equations (PDEs), due to complex geometries, interactions between physical variables, and the lack of large amounts of high-resolution training data. To address these issues, we propose Codomain Attention Neural Operator (CoDA-NO), which tokenizes functions along the codomain or channel space, enabling self-supervised learning or pretraining of multiple PDE systems. Specifically, we extend positional encoding, self-attention, and normalization layers to the function space. CoDA-NO can learn representations of different PDE systems with a single model. We evaluate CoDA-NO’s potential as a backbone for learning multiphysics PDEs over multiple systems by considering few-shot learning settings. On complex downstream tasks with limited data, such as fluid flow simulations and fluid-structure interactions, we found CoDA-NO to outperform existing methods on the few-shot learning task by over $36\%$. The code is available at https://github.com/ashiq24/CoDA-NO.
1 Introduction
Figure 1: CoDA-NO, our proposed architecture, generalizes to novel multi-physics systems. CoDA-NO can be pre-trained on fluid dynamics data, a single physics system governed by the Navier-Stocks equation with variables such as velocities ($u_{x},u_{y}$) and pressure $(p)$. The pre-trained CoDA-NO can be easily adapted to a multi-physics fluid-solid interaction system governed by coupled Navier-Stocks and Elastic wave equations. It handles the addition displacement variables ($d_{x},d_{y}$) without any architecture changes.
Many science and engineering problems frequently involve solving complex multiphysics partial differential equations (PDEs) (strang2007computational). However, traditional numerical methods usually require the PDEs to be discretized on fine grids to capture important physics accurately. This becomes computationally infeasible in many applications.
Neural operators (li2021fourier; azzizadenesheli2023neural) have shown to be a powerful data-driven technique for solving PDEs. Neural operators learn maps between function spaces and converge to a unique operator for any discretization of the functions. This property, called discretization convergence, makes them suitable for approximating solution operators of PDEs. By training on pairs of input and solution functions, we obtain estimates of solution operators that are often orders of magnitude faster to evaluate than traditional PDE solvers while being highly accurate (schneider2017climate; kossaifi2023multigrid; bonev2023sfno).
Neural operators are a data-driven approach and therefore, their performance depends on the quality and abundance of training data. This can be a bottleneck since it is expensive to generate data from traditional solvers or collect them from experiments. A promising approach to dealing with this is to learn physically meaningful representations across multiple systems of PDEs and then transfer them to new problems. Such a self-supervised learning approach with foundation models has found immense success in computer vision and natural language processing (caron2021emerging; radford2021learning). Foundation models are trained in a self-supervised manner on large unlabeled datasets. They can be efficiently adapted or fine-tuned to a broad range of downstream tasks with minimal to no additional data or training.
Recent works have delved into the possibility of establishing a foundation model for solving PDEs (subramanian2023towards; mccabe2023multiple). However, these methods only work on simple, predetermined PDEs with a fixed number of interacting variables. In addition, they are restricted to uniform equidistant grids, which limits their applicability. For example, standard patching-based approaches, used in Vision Transformers (ViTs) (dosovitskiy2020image), often struggle with discontinuities in predicted functions (zhang2022styleswin). Further, since they are limited to fixed uniform grids, they cannot be applied to resolutions different from the training resolution.
Our Approach. We propose a novel co-domain transformer neural operator (CoDA-NO) that alleviates the above problems. While previous neural operators have successfully modeled solution operators of one PDE system, our transformer architecture explicitly models dependencies across different physical variables of multiple PDE systems, using self-attention in the co-domain or channel space. Specifically, CoDA-NO tokenizes the functions channel-wise in each layer and treats each physical variable as a token, obviating the need for patching. It also extends the transformer from finite-dimensional vectors to infinite-dimensional functions by appropriately redesigning positional encodings, self-attention mechanism, and normalization.
CoDA-NO can be applied to varying numbers of input functions (on different geometries) and easily adapt to novel PDEs with additional or fewer interacting variables, as illustrated in Fig. 1. This allows us to learn multiple PDE systems in one model. Our proposed framework is highly modular and allows the inclusion of previous neural operators as components. In contrast to the previous works by li2023scalable; guibas2021adaptive that extend transformers to operator learning by designing attention on the spatial or temporal domains, in CoDA-NO, we instead design attention on the co-domain or channel space to capture the different physical variables present in multiphysics PDEs and across different PDEs with common physical variables (see Fig. 1).
To assess models’ generalizability across diverse physical systems, we examine two problems. One involves fluid dynamics governed by the Navier-Stokes equation, with velocity and pressure as variables. The other is a fluid-structure interaction problem governed by both the Navier-Stokes equation and the Elastic wave equation, with an additional variable of displacement field associated with the Elastic wave equation. Both of the problems also provide an additional challenge of irregular mesh over a complex geometry.
We pre-train CoDA-NO in a self-supervised manner on snapshots of fluid flows by masking different parts of the velocity or pressure fields. Using few-shot supervised fine-tuning, we show that our model can adapt to unseen viscosities and additional displacement fields given by the elastic wave equation. We use graph neural operator (GNO) layers li2023gino as encoders and decoders to handle time-varying irregular meshes of fluid-structure interaction problems. For the few-shot learning problem, our model achieves $36.8\%$ lower errors on average compared to the best-performing baseline trained from scratch on the target problem.
Our contributions are as follows:
  • [leftmargin=*,topsep=0pt, itemsep=0pt]
  • We propose a co-domain attention neural operator that efficiently learns solution operators to PDEs by formulating transformer operations in function space and ensuring discretization convergence.
  • The proposed architecture enables self-supervised learning in function space for diverse physical systems by handling varying numbers of input functions and geometries.
  • CoDA-NO achieves state-of-the-art performance in generalizing to unknown physical systems with very limited data. That is, CoDA-NO can be viewed as the first foundation neural operator for multiphysics problems.
2 Related Works
Transformers for solving PDEs.
Recently, takamoto2023learning proposes a method to weight variables/codomains of the input function based on the weights calculated from the PDE parameters. Another study by li2023scalable proposes a scalable transformer architecture by combining a projection operator to a one-dimensional domain and a learnable factorized kernel. In contrast to these works, CoDA-NO provides a complete attention operator by considering each co-domain as a token function, i.e., infinite dimensional vector, extending traditional transformers on finite dimension tokens.
Self-supervised learning.
Self-supervised learning (SSL) was proposed to tackle the issue of limited labeled data he2022masked; radford2018improving; chen2020simple. It allowed the training of large foundation models on massive amounts of data in the field of computer vision and natural language processing. These models are being successfully applied to a wide range of downstream tasks with minimal to no additional task-specific data (chowdhery2023palm; saharia2022photorealistic; liu2021swin; radford2021learning).
Pre-training for PDE solving.
Self-supervised pre-trained models have also been gaining traction in the domain of scientific computing. mccabe2023multiple propose pretraining the models with autoregressive tasks on a diverse dataset of multiple PDEs. These models can then be fine-tuned for specific downstream PDEs. Several recent studies have investigated task-agnostic approaches through masking-and-reconstruction (he2022masked) and the consistency of representations under symmetry transformations (xu2023self; li2020fourier; mialon2023self). Recent work by subramanian2023towards sheds light on the transferability of these models between different systems of PDEs. While these methods achieve good performance, the target (downstream) PDE must maintain a strict resemblance to the ones used for pretraining. In addition, adapting these models for PDEs with new additional output variables is not possible. ViT-based patching approaches (zhang2022styleswin) disrupt the continuity and are not resolution-invariant.
3 Preliminaries
Figure 2: On the left, we illustrate the architecture of the Codomain Attention Neural Operator. Each physical variable (or co-domain) of the input function is concatenated with variable specific positional encoding (VSPE). Each variable, along with the VSPE, is passed through a GNO layer, which maps from the given non-uniform geometry to a latent regular grid. Then, the output on a uniform grid is passed through a series of CoDA-NO layers. Lastly, the output of the stacked CoDA-NO layers is mapped onto the domain of the output geometry for each query point using another GNO layer. On the right, we illustrate the mechanism of codomain attention. At each CoDA-NO layer, the input function is tokenized codomain-wise, and each token function is passed through the ${\mathcal{K}},{\mathcal{Q}},$ and ${\mathcal{V}}$ operators to get key, query, and value functions $\{k^{1},k^{2}\},\{q^{1},q^{2}\},$ and $\{v^{1},v^{2}\}$ respectively. The output function is calculated via an extension of the self-attention mechanism to the function space.
We briefly review the necessary concepts to understand our approach and establish the notation. For any input function $a\colon\mathcal{D}\rightarrow\mathbb{R}^{d_{in}}$, we will denote the $d_{in}$ dimensional output shape as codomain. We consider each of the components of the codomain as different physical variables, which are real-valued functions over the input domain $\mathcal{D}$, i.e., $a=[a^{1},\dots,a^{d_{in}}]$ with $a^{i}:\mathcal{D}\rightarrow\mathbb{R}$. The same applies to the output function $u\colon\mathcal{D}\rightarrow\mathbb{R}^{d_{out}}$.
Neural Operators. Neural operators are a class of deep learning architectures designed to learn maps between infinite-dimensional function spaces (kovachki2021neural). A Neural Operator seeks to approximate an operator $\mathcal{G}$ that maps an input function $a\in\mathcal{A}$ to its corresponding output function $u\in\mathcal{U}$ by building a parametric map $\mathcal{G}_{\phi}:\mathcal{A}\to\mathcal{U}$. The typical architecture of a Neural Operator can be described as
$\mathcal{G}_{\phi}=\mathcal{P}\circ{\mathcal{I}}_{L}\circ\dots\mathcal{I}_{1}\circ\mathcal{L}.$ (1)
Here, $\mathcal{L}\colon a\rightarrow w_{0}$ and $\mathcal{P}\colon w_{L}\rightarrow u$ are lifting and pointwise projection operators, respectively. The action of any pointwise operator :${\mathcal{H}}:\{f:\mathcal{D}\rightarrow\mathbb{R}^{d_{f}}\}\rightarrow\{g:\mathcal{D}\rightarrow\mathbb{R}^{d_{g}}\}$ can be defined as
${\mathcal{H}}[f][x]=h_{\phi}(f(x)),$ (2)
where $h_{\phi}:\mathbb{R}^{d_{f}}\rightarrow\mathbb{R}^{d_{g}}$ is any function with parameters $\phi$. The integral operator ${\mathcal{I}}_{l}:w_{l-1}\rightarrow w_{l}$ performs a kernel integration over the input function $w_{l-1}$ as
${\mathcal{I}}_{l}[w_{l-1}](x)=\int_{\mathcal{D}_{l-1}}k_{l}(x,y)w_{l-1}(y)\,\mathrm{d}y.$ (3)
Here, $\mathcal{D}_{l-1}$ is the domain of the function $w_{l-1}$. In the case of Fourier Neural operators (FNO) (li2020fourier), a convolution kernel, i.e., $k_{l}(x,y)=k_{l}(x-y)$ was used. By the convolution theorem, this enables the representation of an integral operator as a pointwise multiplication of the Fourier coefficients as follows:
$w_{l}=\mathcal{F}^{-1}(\mathcal{F}(k_{l})\odot\mathcal{F}(w_{l-1})).$ (4)
In the presence of discretization, the corresponding integration is approximated using a discrete Fourier transform,  i.e., the Riemann sum approximation, and for equidistant grids, can be efficiently computed using a fast Fourier transform (FFT) implementation.
For the Graph neural operator (GNO) (li2020neural), a small neighborhood $B_{r}(x)\cap\mathcal{D}_{l-1}$ around the point $x$ is considered instead of integrating over the whole domain $\mathcal{D}_{l-1}$, such that Eq. (3) changes to
$w_{l}(x)=\int_{B_{r}(x)\cap D_{l-1}}k_{l}(x,y)w_{l-1}(y)\,\mathrm{d}y.$ (5)
Given a set of evaluations of the function $w_{l-1}$ on points $\{y_{i}\}_{i=1}^{n}\subset\mathcal{D}_{l-1}$, the kernel integral can be approximated by
$w_{l}(x)\approx\sum_{y_{i}\in B_{r}(x)}k_{l}(x,y_{i})w_{l-1}(y_{i})q_{i},$ (6)
where $q_{i}\in\mathbb{R}$ are suitable quadrature weights (kovachki2021neural). The discretized kernel integral can be viewed as a message passing on graphs, where the neighborhood of each point $x$ consists of all points within radius $r$.
4 Method
Problem Statement. Our goal is to develop a neural operator architecture $\mathcal{G}$ that explicitly models the interaction between physical variables of PDE systems. To learn and predict different systems, the architecture should not be restricted to a fixed number of such variables.
Let’s consider two input functions $a\colon\mathcal{D}\rightarrow\mathbb{R}^{d_{in}}$ and $\tilde{a}\colon\mathcal{D}\rightarrow\mathbb{R}^{\tilde{d}_{in}}$ of two different PDE solution operators with corresponding output functions $u\colon\mathcal{D}\rightarrow\mathbb{R}^{d_{out}}$ and $\tilde{u}\colon\mathcal{D}\rightarrow\mathbb{R}^{\tilde{d}_{out}}$. In general, the functions $a$ and $\tilde{a}$ represent $d_{in}$ and $\tilde{d}_{in}$ physical variables over the domain $\mathcal{D}$ with $d_{in}\neq\tilde{d}_{in}$ We aim to design neural operator architectures $\mathcal{G}$ that can both be applied to $a$ as well as $\tilde{a}$ despite the different codomains of the input as well as output functions. Such property provides the possibility to evaluate or finetune the operator on PDEs with different numbers of variables than those on which it was trained. In particular, when the PDE systems have overlapping physical variables $\{a^{i}\}_{i=1}^{d_{in}}\cap\{\tilde{a}^{i}\}_{i=1}^{\tilde{d}_{in}}\neq\emptyset$, this naturally allows to transfer learned knowledge from one system to the other. We will next describe the details of the CoDa-NO layers and architecture to achieve this goal.
Permutation Equivariant Neural Operator. As we consider the vector-valued input function $a$ as a set of $d_{in}$ functions $\{a^{1},a^{2},\dots,a^{d_{in}}\}$ that represent different physical variables of the PDE. We seek to construct operators that act on sets of input functions with different cardinalities. For an efficient implementation, we mimic transformer architectures and share weights across different variables. We achieve this by defining permutation equivariant integral operator ${\mathcal{I}}_{per}$ as
${\mathcal{I}}_{per}[w]=\bigg{[}{\mathcal{I}}[w^{1}_{e}],\dots,{\mathcal{I}}[w^{d_{in}}_{e}]\bigg{]},$ (7)
[…]
CoDA-NO Layer. Given a function $w:\mathcal{D}\rightarrow\mathbb{R}^{d}$, we partition the function into a set of so-called token functions $w^{i}:\mathcal{D}\rightarrow\mathbb{R}^{d^{\prime}}$ for $i\in\{1,\dots L\}$ along the codomain, such that $w=\big{[}w^{1},\dots w^{T}\big{]}$. That is, $w$ represents the codomain-wise concatenation of the token functions $w_{i}$ and $d^{\prime}=\frac{d}{L}$. If no other value is specified, we assume that $d^{\prime}=1$. The CoDA-NO layer now processes the token functions using an extension of the self-attention mechanism to the function space (see Appendix Sec. LABEL:app:attention and Fig. 2).
Let us begin by introducing a single-head CoDA-NO layer. Later, we will expand the concept to multi-head codomain attention. We extend the key, query, and value matrices of the standard attention (see Appendix Sec. LABEL:app:attention for details) to operators mapping token functions $w^{i}\colon\mathcal{D}\to\mathbb{R}^{d^{\prime}}$ to key, query, and value functions. We define the key, query, and value operators as
${\mathcal{K}}:w^{i}\rightarrow\{k^{i}:\mathcal{D}\rightarrow\mathbb{R}^{d_{k}}\},$ (8)
${\mathcal{Q}}:w^{i}\rightarrow\{q^{i}:\mathcal{D}\rightarrow\mathbb{R}^{d_{q}}\},$ (9)
${\mathcal{V}}:w^{i}\rightarrow\{v^{i}:\mathcal{D}\rightarrow\mathbb{R}^{d_{v}}\}.$ (10)
Assuming $d_{k}=d_{q}$, we denote by $k^{i}={\mathcal{K}}[w^{i}]$, $q^{i}={\mathcal{Q}}[w^{i}]$, and $v^{i}={\mathcal{V}}[w_{i}]$ the key, query, and value functions of the token functions, respectively.
Next, we calculate the output (token) functions $o^{i}:\mathcal{D}\rightarrow\mathbb{R}^{d_{v}}$ as
$o^{i}={{\tt Softmax}}\left([\frac{\langle q^{i},k^{1}\rangle}{\tau}\dots\frac{\langle q^{i},k^{T}\rangle}{\tau}]\right)[v^{1},\dots,v^{L}]^{\top}$ (11)
where $\tau$ is the temperature hyperparameter. Here, .,$\langle$.,$\rangle$ . , . ⟩ denotes a suitable dot product in the function space. We take the $L^{2}(\mathcal{D},\mathbb{R}^{d_{k}})$-dot product given by
$\langle q^{i},k^{j}\rangle=\int_{\mathcal{D}}\langle q^{i}(x),k^{j}(x)\rangle\,\mathrm{d}x,$ (12)
where the integral can be discretized using quadrature rules, similar to the integral operator (see Sec. 3).
To implement multi-head attention, we apply the (single-head) attention mechanism described above separately for multiple heads $h\in\{1,\dots H\}$ using ${\mathcal{K}}^{h},{\mathcal{Q}}^{h},\text{ and }{\mathcal{V}}^{h}$ to obtain $o^{i,h}$. We then concatenate these outputs $o^{i,h}$ along the codomain and get $c^{i}:=[o^{i,1},\dots o^{i,H}]$. Finally, we use an operator
:$\mathcal{M}:\{c^{i}:\mathcal{D}\rightarrow\mathbb{R}^{H\cdot d_{v}}\}\rightarrow\{o^{i}:\mathcal{D}\rightarrow\mathbb{R}^{d_{v}}\}$ (13)
to get the output function $o^{i}$.
We obtain the output of the attention mechanism by concatenating $o^{i}$ as $o=[o^{1},o^{2},\dots o^{L}]$. Finally, we complete the CoDA-NO layer by applying a permutation equivariant integral operator ${\mathcal{I}}_{per}$ on $o$. When CoDA-NO is acting on functions sampled on a uniform grid, the internal operators $\mathcal{K}^{h},{\mathcal{Q}}^{h},\mathcal{V}^{h},\mathcal{M},$ and ${\mathcal{I}}$ are implemented as FNOs.
Function Space Normalization. Normalization is a vital aspect of deep learning architectures. However, when it comes to neural operators mapping infinite-dimensional functions, this topic remains largely unexplored. We now provide a natural extension. Given a function $w$, let $w^{j}_{e}:\mathcal{D}\rightarrow\mathbb{R}^{d_{l}}$ be the codomain group. Then we calculate the mean $\mu\in\mathbb{R}^{d_{l}}$ and standard deviation $\sigma\in\mathbb{R}^{d_{l}}$ for each codomain group as
$\mu^{j}=\int_{\mathcal{D}}w^{j}_{e}(x)\,\mathrm{d}x.$ (14)
$\sigma^{j}=\bigg{(}\int_{\mathcal{D}}(w^{j}_{e}(x)-\mu^{j})^{\circ 2}\,\mathrm{d}x\bigg{)}^{\circ\frac{1}{2}}.$ (15)
Here, ${\circ i}$ denotes the elementwise (Hadamard) $i^{th}$-power. The normalization operator can be written as
${\tt Norm}[w^{j}_{e}](x)=({\mathbf{g}}\oslash\sigma^{j})\odot(w^{j}_{e}(x)-\mu^{j})+{\mathbf{b}},$ (16)
[…]
Variable Specific Positional Encoding (VSPE). We learn positional encoders $e^{j}:\mathcal{D}\rightarrow\mathbb{R}^{d_{en}}$ for each physical variable $j\in\{1,\dots,d_{in}\}$, for the given vector-valued input function [^1] For ease of presentation, we assume all variable $a_{i}$ are defined on the same domain $\mathcal{D}$, but our method can easily be extended to accommodate different domains. $a=[a^{1},\dots,a^{d_{in}}]$. We concatenate each positional encoding $e^{j}$ with the respective variable $a^{j}:\mathcal{D}\rightarrow\mathbb{R}$ along to codomain to obtain extended input functions $a^{j}_{e}=[a^{j},e^{j}]$. Next, we apply a shared pointwise lifting operator
$\mathcal{P}:\{a_{e}^{j}:\mathcal{D}\rightarrow\mathbb{R}^{d_{en}+1}\}\rightarrow\{w^{j}_{e}:\mathcal{D}\rightarrow\mathbb{R}^{d}\},$ (17)
typically with $d>d_{en}+1$. Finally, we concatenate $w^{j}_{e}$, $j\in\{1,\dots d_{in}\}$, to get the lifted function
$w=[w^{1}_{e},\dots,w^{d_{in}}_{e}]\colon\mathcal{D}\to\mathbb{R}^{d\cdot d_{in}}.$ (18)
We refer to each $w^{j}_{e}$ as a codomain group.
[…]
Figure 3: Test time adaptation to new physical variables. The model is pre-trained on the Navier-Stokes equation dataset, which comprises physical variables such as velocities $u_{x}$, $u_{y}$, and pressure $p$. To adapt this pre-trained model on a fluid-solid interaction dataset containing an additional Elastic wave equation with new displacement variables $d_{x}$ and $d_{y}$, it is only necessary to add two additional VSPEs to the whole pipeline.
Figure 4: Horizontal velocity $u_{x}$ at $t$ and $t+\delta t$ time step.
Model Training. The model undergoes a two-stage training process: Self-supervised pretraining is followed by a supervised fine-tuning stage. For a summary, we refer to Fig. 3.
Pre-training. In the context of self-supervised pretraining, the objective is to train the model to reconstruct the original input function from its masked version. Within this phase, the model’s encoding component is denoted as the Encoder, while the decoding component comprises the Reconstructor. The values of the input function at a specific percentage of mesh points are randomly masked to zero, and certain variables (of the co-domain) of the input function are entirely masked to zero. The model is then trained to reconstruct the original input from this masked version. We emphasize that the self-supervised learning phase is agnostic of the downstream supervised task and only requires snapshots of simulations of the physical systems.
Fine-tuning. In the supervised fine-tuning phase, the Reconstructor is omitted from the decoding module and replaced by a randomly initialized Predictor module. The parameters of the Encoder and VSPEs are copied from pre-trained weights. If the fine-tuning (target) PDE introduces variables that are not present in the pre-training PDE, we train additional variable encoders for these newly introduced variables (see Fig. 3). This ensures that the model adapts to the expanded set of variables needed for the fine-tuning task.
5 Experiments
Modeling Fluid-structure Interaction. We consider the following problems: (a) a fluid dynamics problem, where a Newtonian, incompressible fluid impinges on a rigid object, and (b) a fluid-structure interaction problem between a Newtonian, incompressible fluid and an elastic, compressible solid object turek2006proposal. We denote $\Omega^{f}_{t}$ (resp. $\Omega^{s}_{t}$) as the domain occupied by the fluid (resp. the solid) at time $t$. The dynamics of the fluid are governed by the Navier-Stokes equations
$&\rho^{f}\frac{\partial u}{\partial t}+\rho^{f}\nabla\cdot(u\otimes u)=\nabla\cdot\bm{\sigma}^{f},~{}\nabla\cdot u=0,\\ \end{aligned}\quad\mbox{in }\Omega^{f}_{t}$ (19)
where $u$ and $\rho^{f}$ denote the fluid velocity and density, respectively. And $\bm{\sigma}^{f}$ denotes the Cauchy stress tensor, given by
$\bm{\sigma}^{f}=-p\mathbb{I}+\mu(\nabla u+\nabla u^{T}),$ (20)
where $\mathbb{I}$ is the identity tensor, $p$ the fluid pressure, and $\mu$ the fluid dynamic viscosity.
For fluid-structure interaction, the deformable solid is governed by the elastodynamics equations
$\rho^{s}\frac{\partial^{2}d}{\partial t^{2}}=\nabla.(J\bm{\sigma}^{s}(\mathbf{F}^{-1})^{T})\qquad\mbox{ in }\Omega^{s}_{t}$ (21)
with $\mathbf{F}=\mathbb{I}+\nabla d$ and $J=\det(\mathbf{F})$. Here $d$, $\rho^{s}$, $F$, and $\bm{\sigma}^{s}$ denote the deformation field, the solid density, the deformation gradient tensor, and the Cauchy stress tensor, respectively(see Eq. (LABEL:solid_stress_tensor) in the Appendix).
The fluid dynamics (resp. the fluid-structure interaction) problem considers a fluid flow past a fixed, rigid cylinder, with a rigid (resp. elastic) strap attached; details regarding the geometric setup (see Fig. 4), as well as the initial conditions, are provided in the Appendix (see Sec. LABEL:app:dataset).
Time-dependent inlet boundary conditions consist of $4^{th}$ order polynomials velocity profiles which vanish at the channel walls (kachuma2007linear; ghosh2017relative). The inlet conditions are given by
$u^{\mathcal{I}}_{c_{1},c_{2}}(y,t)=v(t)\cdot\frac{y(y-H)\left(y-c_{1}\frac{H}{2}\right)\left(y-c_{2}\frac{H}{2}\right)}{H(1-c_{1})(1-c_{2})}.$ (22)
Here $v$ is the ramp function defined as
$v(t)=\left\{\begin{array}[]{ccl}70\cdot\left(1-\cos\left(\frac{\pi t}{2}\right)\right)&\mbox{if}&0\leq t<2\\ 140&\mbox{if}&t\geq 2\\ \end{array}\right.$ (23)
and $(c_{1},c_{2})\in\mathcal{I}$, where
=$\mathcal{I}=\left\{(a,b)\in\{-6,-4,-2,0,2,4,6\}^{2}\ |\ a\leq b\right\}$ (24)
are enforced at the inlet $x=0$. Details regarding the boundary conditions imposed on the strap, which forms the solid-fluid interface, as well as on the remaining domain boundaries, are provided in the Appendix (see Sec. LABEL:app:dataset).
Dataset Description and Generation. Two datasets are generated using the TurtleFSI package, one for fluid-structure interaction and the other for fluid dynamics (bergersen2020turtlefsi). We simulate the fluid-structure interaction and the fluid dynamics test cases described above up to time $T=10$, using a constant time-step $\delta t=\frac{T}{n}$, where $n=1000$. The data sets are composed of solution trajectories $[u_{t},p_{t},d_{t}]$ (resp. $[u_{t},p_{t}]$), which denote the approximate solution of the fluid-structure interaction problem (resp. the fluid dynamics problem) at times $t=i\delta t,i\in\{0,\dots,n\}$. These trajectories are obtained on the basis of all triplets $(\mu,c_{1},c_{2})$ describing combinations of fluid viscosities and inlet conditions, such that $\mu\in\{1,5,10\},(c_{1},c_{2})\in\mathcal{I}.$ Additional details regarding the solver and simulation parameters used can be found in the Appendix (see Sec. LABEL:app:dataset).
Table 1: Test errors ($L_{2}$ loss) for fluid dynamics (NS) and fluid-solid interaction (NS+EW) datasets with viscosity $\mu=5.0$ for different numbers of few-shot training samples.
[…]
Pretrain
Dataset
# Few Shot Training Samples
5 25 100
NS NS+EW NS NS+EW NS NS+EW
GINO - 0.200 0.122 0.047 0.053 0.022 0.043
DeepO - 0.686 0.482 0.259 0.198 0.107 0.107
GNN - 0.038 0.045 0.008 0.009 0.008 0.009
[…]
NS+EW 0.024 0.040 0.006 0.005 0.005 0.003
Experiment Setup. To evaluate the adaptability of the pre-trained model to new PDEs featuring varying numbers of variables, we conduct two distinct pre-training procedures for CoDA-NO and obtain two pre-trained models: ${\tt\mathcal{G}^{p}_{NS-EW}}$ and ${\tt\mathcal{G}^{p}_{NS}}$. The former is pre-trained on a fluid-structure interaction dataset that combines the Navier-Stokes equation and the elastic wave equation, denoted as ${\tt\mathcal{G}^{p}_{NS-EW}}$. The latter, ${\tt\mathcal{G}^{p}_{NS}}$, is pre-trained on a dataset containing fluid motion governed solely by the Navier-Stokes equation. In both scenarios, the pre-training involves utilizing 8000 snapshots of flow and displacement fields, encompassing fluid viscosities $\mu\in\{1,10\}$.
The supervised task involves training the model to predict the state of the system at the subsequent time step based on its current state. For the fluid-structure interaction dataset, we train an operator ${\tt\mathcal{G}_{NS-EW}}$ such that
${\tt\mathcal{G}_{NS-EW}}:[u_{t},p_{t},d_{t}]\rightarrow[u_{t+\delta t},p_{t+\delta t},d_{t+\delta t}],$ (25)
where $u,p$, and $d$ are the velocity, pressure, and mesh deformation fields (see Sec. 5). For the data with only fluid motion, we train the operator ${\tt\mathcal{G}_{NS}}$ which maps between the current and next time step velocity and pressure field as
${\tt\mathcal{G}_{NS}}:[u_{t},p_{t}]\rightarrow[u_{t+\delta t},p_{t+\delta t}].$ (26)
For both datasets, the pre-trained model is fine-tuned for unseen viscosity $\mu=5.0$ with different numbers of few shot examples. The inlet conditions of these simulations are excluded from the pre-training data. So both viscosity and inlet conditions of the target PDEs are absent in the per-taining dataset. We finetune both of the pre-train models ${\tt\mathcal{G}^{p}_{NS-EW}}$ and ${\tt\mathcal{G}^{p}_{NS}}$ on each dataset.
Baselines. For comparision on the supervised tasks, we train GINO (li2023gino), DeepONet (lu2019deeponet), and a graph neural network (GNN) from scratch following battaglia2016interaction. To efficiently handle irregular mesh, in the branch network of DeepONet we use a GNN layer followed by MLPs. It should be noted that employing the existing models for pertaining and subsequent finetuning on the target datasets is nontrivial due to complex geometry and the changes in the number of physical variables between the pertaining and target datasets. We report the $L^{2}$ error between the predicted and target functions, which serves as a measure of model performance .
Figure 5: Test Error of different models for different number of training samples.
Implementation Details. The variable encoders are implemented using a multi-layer perceptron mapping the position $x\in\mathcal{D}$ to the embedding vector. Following transformers and NeRF (mildenhall2021nerf) model, we use positional encoding instead of raw coordinates. The Encoder and Reconstructor modules use three stacked CoDA-NO layers. The Predictor modules uses one layer of CoDA-NO.
For masking, 50% of the mesh points of 60% of variables are masked, i.e., set to zero, or 30% of the variables are masked out completely (see Appendix Sec. LABEL:app:masking). For every training sample, one of these two masking choices is selected with equal probability.
Results. In Tab. 1, we report the $L^{2}$ error between the predicted and target functions for each of the methods. We observe that the pre-trained CoDA-NO model performs better than the baselines trained from scratch. Importantly, the performance gain is higher when the number of few-shot examples is very low. This demonstrates the sample efficiency and generalization capability of CoDA-NO to previously unseen physical systems.
Next, when CoDA-NO is pre-trained solely on the Navier-Stokes equation (NS dataset), it shows an impressive ability to adapt to the more challenging fluid-structure interaction PDE (NS+EW dataset). Finally, when CoDA-NO is pre-trained on the more intricate NS+EW dataset, it easily adapts to the simpler NS dataset through fine-tuning.
Table 2: Ablation studies on fluid dynamics (NS) and fluid-solid interaction (NS+EW) datasets with viscosity $\mu=5.0$ for different numbers of few-shot training samples; Reporting $L_{2}$ loss.
[…]
Pretrain
Dataset
# Few Shot Training Samples
5 25 100
NS NS+EW NS NS+EW NS NS+EW
ViT - 0.271 0.211 0.061 0.113 0.017 0.020
U-Net - 13.33 3.579 0.565 0.842 0.141 0.203
Ours - 0.182 0.051 0.008 0.084 0.006 0.004
[…]
NS+EW 0.024 0.040 0.006 0.005 0.005 0.003
Ablation Studies. To evaluate the efficiency of the proposed CoDA-NO layer, VSPE, and codomain-wise tokenization, we compare it against the vision transformer (ViT) (dosovitskiy2020image) and the Unet (ronneberger2015u) model.
The mesh points of the NS and NS+EW datasets are irregular and change for each sample. As a result, we cannot directly train ViT and Unet on the dataset. Instead, we follow the architecture of GINO and use a GNN layer to query the latent function on a uniform grid and apply Unet and ViT on the uniform grid, followed by another GNN layer to get the output at the desired query points. We report these comparisons in Tab. 2. On few-shot learning tasks, pre-trained CoDA-NO models outperform the regular transformer base ViT, U-Net models, and CoDA-NO trained from scratch.
We also assess the significance of pre-training, we add a CoDA-NO baseline trained from scratch for each of the datasets. Specifically, we evaluate all the models on all viscosities $\mu\in\{1.0,5.0,10.0\}$ with different numbers of training samples where the initial condition used in these simulations is excluded from the pre-training dataset. The ablation results are shown in Fig. 5.
We observe that when the number of supervised training data is low, the pre-trained CoDA-NO model outperforms other models on most of the dataset. Conversely, with more training samples, the CoDA-NO trained from scratch achieves similar or better performance than the pre-trained model.
Interestingly, when more training data is available, CoDA-NO trained from scratch significantly outperforms all other models, including the regular attention mechanism (ViT), U-Net, and GNN. This illustrates the applicability of the CoDA-NO architecture not just for few-shot learning but also in scenarios where abundant data is available, thanks to its effective attention mechanism for function spaces. We also provide an ablation study where we freeze the parameters of the "Encoder" and only train the parameters of the "Predictor". Here we also perform significantly better than other models on most of the dataset (see Appendix Sec. LABEL:app:ablation_forzen).
Author Contributions
If you’d like to, you may include a section for author contributions as is done in many journals. This is optional and at the discretion of the authors.
Acknowledgments
Use unnumbered third level headings for the acknowledgments. All acknowledgments, including those to funding agencies, go at the end of the paper.
Appendix A Appendix
You may include other additional sections here.